OpenStudy (anonymous):

Find the remainder when f(x) is divided by (x - k) f(x) = 4x^3 - 5x^2+2x+11; k= 3

OpenStudy (shubhamsrg):

f(x) = g(x) (x-k) + r f(k) = 0 + r was that any help ?

OpenStudy (anonymous):

(x-k) = (x-3)

OpenStudy (anonymous):

Remainder theorem f(3)=4x^3 - 5x^2+2x+11

OpenStudy (anonymous):

That didn't help.

OpenStudy (anonymous):

put x = 3 in the eq and solve

OpenStudy (anonymous):

Oh, okay!

OpenStudy (anonymous):

Thanks.